OpenStudy (anonymous):

The twice–differentiable function f is defined for all real numbers and satisfies the following conditions: f(0)=3 f′(0)=5 f″(0)=7 The function h is given by h(x)=cos(kx)[f(x)]+sin(x) for all real numbers, where k is a constant. Find h ′(x) and write an equation for the line tangent to the graph of h at x=0.

OpenStudy (anonymous):

@ganeshie8 please help!!

OpenStudy (anonymous):

will give medal :) @Loser66 @ganeshie8

OpenStudy (loser66):

for h'(x) , just take derivative h(x) for tangent, replace x =0 in h'(x), what do you have?

OpenStudy (anonymous):

How do i take the derivative of h(x)

OpenStudy (loser66):

what??? really?? you don't know how to take derivative of a function?

OpenStudy (anonymous):

no :( not this function this ones complicated

OpenStudy (anonymous):

cos(kx)[f(x)]+sin(x)

OpenStudy (loser66):

ok, (sin (x))' = ?

OpenStudy (anonymous):

idk :(

OpenStudy (loser66):

@jim_thompson5910 what should I do?

jimthompson5910 (jim_thompson5910):

what is the derivative of cosine?

OpenStudy (anonymous):

im not sure ive been slacking so badly

jimthompson5910 (jim_thompson5910):

you should have something like this in your notes http://sub.allaboutcircuits.com/images/11045.png

OpenStudy (loser66):

@1019.jams We can help you cook only when you have something to cook. If you have nothing, how can we help?

OpenStudy (anonymous):

can you please show all the steps then i'll understand from there :) that's how i learn math i always follow the steps that my teachers show me

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

@satellite73

jimthompson5910 (jim_thompson5910):

1019.jams, did you look at the link I posted

OpenStudy (anonymous):

yes; now i need to write an equation for the line tangent to the graph of h at x=0 @Loser66

OpenStudy (anonymous):

@ganeshie8

jimthompson5910 (jim_thompson5910):

so if y = sin(x), then what is dy/dx equal to?

OpenStudy (anonymous):

cos(x) :)

jimthompson5910 (jim_thompson5910):

if y = cos(x), then dy/dx = ???

OpenStudy (anonymous):

this is what i have so far h'(x) = [cos (x)(f(x)]' +(sin (x))'= (cos(x))' f(x) + cos(x) f'(x) + cos (x) = sin x f(x) +cos x f'(x) +cos x

OpenStudy (anonymous):

if you are stuck at taking derivatives, this problem is not for you

jimthompson5910 (jim_thompson5910):

don't forget that you had cos(kx)*f(x) and not just cos(x)*f(x)

jimthompson5910 (jim_thompson5910):

so you're missing that k

OpenStudy (anonymous):

got it :) what do i do next :)

OpenStudy (anonymous):

how do i find an equation for the line tangent to the graph of h at x=0

jimthompson5910 (jim_thompson5910):

so what is h'(x) equal to

OpenStudy (anonymous):

h'(x) = [cos (kx)(f(x)]' +(sin (x))'= (cos(kx))' f(x) + cos(kx) f'(x) + cos (kx) = sin x f(x) +cos kx f'(x) +cos kx

jimthompson5910 (jim_thompson5910):

incorrect

jimthompson5910 (jim_thompson5910):

look back at the original problem, and be careful about the chain rule